OpenStudy (loser66):

Let a, b be integers st (3,a) = (3, b) =1. Show that a^2 + b^2 is not the square of an integer. Please, help

jimthompson5910 (jim_thompson5910):

(3,a) = 1 means that the GCD of 3 and a is 1 So a does not equal 3 or any of its multiples which means a = 3k+1 or a = 3k+2 Similar logic applies to b as well since (3,b) = 1 too. ------------------------------------------------ You'll need to break things down into 3 cases Case 1: a = 3k+1, b = 3k+1 Case 2: a = 3k+1, b = 3k+2 Case 3: a = 3k+2, b = 3k+2 ------------------------------------------------ Proving Case 1 a = 3k+1 b = 3k+1 k is any integer a^2 + b^2 = (3k+1)^2 + (3k+1)^2 a^2 + b^2 = 2*(3k+1)^2 2*(3k+1)^2 is NOT in the form m^2, where m is an integer. Why not? The portion (3k+1)^2 is certainly a perfect square but the 2 out front is not a perfect square Put another way m^2 = 2*(3k+1)^2 m = sqrt(2*(3k+1)^2) m = sqrt(2)*sqrt((3k+1)^2) m = sqrt(2)*(3k+1) So m is irrational and not an integer. This proves a^2 + b^2 is not a perfect square if a = 3k+1, b = 3k+1, where k is any integer

OpenStudy (empty):

Alternate solution: \[a=3n \pm 1 \]\[b=3m \pm 1\] \[a^2 + b^2 \equiv c^2 \mod 3 \]\[(\pm 1)^2 + (\pm 1)^2 \equiv c^2 \mod 3\]\[2 \equiv c^2 \mod 3\] But in mod 3, \(0^2=0\), \(1^2=1\) and \(2^2=1\), so contradiction.

jimthompson5910 (jim_thompson5910):

I'm not thinking. I should have used different integers for a and b like @Empty did. My bad

OpenStudy (empty):

Yeah, I just wanted to have some fun with this, my first instinct was to do the exact same thing you did, but I wanted to find a better way!

OpenStudy (loser66):

I got @jim_thompson5910 method @Empty is it the same, but combine the cases, right? it doesn't derive from (a, 3) =1, right?

OpenStudy (empty):

Yeah this is the same thing as what he did, just in a more convenient notation that allows us to do all the cases at one time because \(a=3n \pm 1\) is really the same as \(3n+1\) and \(3n+2\). Does that answer your question? I don't completely understand your question

OpenStudy (loser66):

Got it. Thanks a lot. It is not my problem. Since my second midterm is upcoming, I search on internet problems to practice. It doesn't have solution. Hence when I am not sure about the solution, I put it here to learn. http://math.scu.edu/~eschaefe/NTpp1.pdf Can you take a look at problem 11? Please

OpenStudy (loser66):

To me, we solve 27x - my = 5 gcd (27, m) = d and d | 5 iff d =1 or =5, but 5 not | 27, hence d =1

OpenStudy (empty):

I think you should remove the x from your equation and just solve \(27=5+my\)

OpenStudy (loser66):

but then, how to solve for m?

OpenStudy (empty):

I have to go eat dinner, I'll be back later, but I would just look at the prime factorizations. There might be more solutions than just this, I need to think a little more about this but you should be able to see that m=2 and m=11 will work I think.

OpenStudy (loser66):

if (27,m) =1, and 27 = 3^3 , hence residue {1,2......26} without multiple of 3. right?

jimthompson5910 (jim_thompson5910):

a = b (mod m) means m | (a-b) 27 = 5 (mod m) means that m | (27-5) ---> m | 22 so m*k = 22 meaning that m is a factor of 22 (k is an integer) so if m = 2, m = 11, or m = 22, then 27 = 5 (mod m) m = 2: 27 = 5 (mod m) 27 = 5 (mod 2) 1 = 1 (mod 2) m = 11: 27 = 5 (mod m) 27 = 5 (mod 11) 5 = 5 (mod 11) m = 22: 27 = 5 (mod m) 27 = 5 (mod 22) 5 = 5 (mod 22) I think those are all the possible values of m but I'm not 100% sure

OpenStudy (loser66):

Wow!! it is so simple. Thank you Mr. Jim

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (loser66):

Mr. Jim, how to do 13? Please